determine the ratio of substance needed to prepare a carbonate buffer of pH= 10.50 providing the available substances are H2CO3, NaHCO3, and Na2CO3. given these equilibrium systems: use Henderson-Hasselbach equation for pH.

HCO^3- <-----> H^+ + CO3^2-
Ka2 =4.8x10^-11 [H^+]???
Answer #1

pKa2    = -log Ka2

           = -log (4.8 x 10^-11)

          = 10.32

pH = pKa2 + log [CO3^2- / HCO3-]

10.50 = 10.32 + log [CO3^2- / HCO3-]

log [CO3^2- / HCO3-] = 0.18

[CO3^2- / HCO3-]   = 1.51

ratio of CO3-2 to HCO3- = [Na2CO3 / NaHCO3]   = 1.51

[H+] = 10^-pH = 10^-10.50

[H+] = 3.16 x 10^-11 M